OpenStudy (anonymous):

help! What are the amplitude, period, and phase shift of the given function? f(t)=2/3cos(4t+3pi)

OpenStudy (anonymous):

f(t)=2/3cos(4t+3pi) standard form y = Acos(Bt +C) =,A=amplitude, Period=2pi/B, Phase shift=C/B

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anytime :)

OpenStudy (aum):

@sangya21 Phase shift = -C/B if y = Acos(Bt + C)

OpenStudy (anonymous):

sorry for missing sign

OpenStudy (dan815):

just look at my beautiful form

OpenStudy (aum):

@dan815 Are you sure S is the shift? In y = Acos(Bt + C), Bt + C = 0 when t = -C/B and that is the phase shift. In y = Acos(2pi/P * x - S), 2pi/P * x - S = 0 when x = SP/(2pi) and that should be the phase shift.

OpenStudy (dan815):

yes ofc

OpenStudy (dan815):

what u said is samething p/2pi = 1/B, and usually we think of right as positive so its -S

OpenStudy (dan815):

@aum except my form is nice because it has A and P and S in the equation itself looks naaiice right

OpenStudy (dan815):

oh isee what u are saying okay

OpenStudy (dan815):

lemme rearrange in shift form

OpenStudy (dan815):